OpenStudy (anonymous):

Please help 2x^-8 is this correct??? 1/2x^8

OpenStudy (anonymous):

definitely...

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

depends. If you mean \[2x^{-8}\] then that is equal to \[\frac{2}{x^8}\] If you mean \[(2x)^{-8}\] then that is equal to \[\frac{1}{(2x)^8}\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

1/256x

OpenStudy (anonymous):

The first one with out ( ) is what I meant ............

OpenStudy (anonymous):

then the answer is \[\frac{2}{x^8}\]

OpenStudy (anonymous):

Okay?! I am a little confused..........

OpenStudy (anonymous):

\[x^{-8}=\frac{1}{x^8}\] So \[2x^{-8} = 2(x^{-8})=2\Big(\frac{1}{x^8}\Big)=\frac{2}{x^8}\]

OpenStudy (anonymous):

okay thanks for breaking it down for me ! :)